Bob fenced in a rectangular garden in his yard. The length of the rectangle is 6 feet longer than the width, and the perimeter is 100 feet. What is the width of the rectangle?
Answer by rfer(16322)

2x+2(x+6)=100
2x+2x+12=100
4x=88
x=22
x+6=28